Delph Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year: Molding Fabrication Total Machine-hours 32,000 42,000 74,000 Fixed manufacturing overhead costs $ 770,000 $ 270,000 $ 1,040,000 Variable manufacturing overhead cost per machine-hour $ 5.40 $ 5.40 During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs: Job D-70: Molding Fabrication Total Direct materials cost $ 376,000 $ 329,000 $ 705,000 Direct labor cost $ 240,000 $ 140,000 $ 380,000 Machine-hours 23,000 9,000 32,000 Job C-200: Molding Fabrication Total Direct materials cost $ 210,000 $ 290,000 $ 500,000 Direct labor cost $ 120,000 $ 280,000 $ 400,000 Machine-hours 9,000 33,000 42,000 Delph had no underapplied or overapplied manufacturing overhead during the year. Exercise 2-15 Part 2 2. Assume Delph uses departmental predetermined overhead rates based on machine-hours. a. Compute the departmental predetermined overhead rates. b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200. c. If Delph establishes bid prices that are 120% of total manufacturing costs, what bid prices would it have established for Job D-70 and Job C-200? d. What is Delph’s cost of goods sold for the year?
Answer #1

Answers

  • All working forms part of the answer
  • Requirement ‘a’

Working

Molding Department:

Fabrication Department:

A

Total Fixed Overheads

$              770,000

$             270,000

B

Total machine hours

                    32,000

                  42,000

C = A/B

Fixed overhead rate

$                   24.06

$                    6.43

D

Variable Overhead rate

$                     5.40

$                    5.40

E = C+D

Plantwide overhead rate = Answer

$                   29.46 per machine hour

$                 11.83 per machine hour

  • Requirement ‘b’

Job D 70

Job C 200

Molding Department:

Direct material cost

$              376,000

$             329,000

Direct Labor cost

$              240,000

$             140,000

Manufacturing Overhead [machine hours x $29.46]

$              677,580

$             265,140

Fabrication Department:

Direct material cost

$              210,000

$             290,000

Direct Labor cost

$              120,000

$             280,000

Manufacturing Overhead [machine hours x $11.83]

$              106,470

$             390,390

Total manufacturing cost assigned = Answer

$          1,730,050

$         1,694,530

  • Requirement ‘c’

Job D 70

Job C 200

A

Total manufacturing cost assigned

$        1,730,050

$               1,694,530

B = A x 120%

Bid Price = Answer

$        2,076,060

$               2,033,436

  • Requirement ‘d’

--Since both the jobs were only jobs completed and sold, the cost of goods sold = total manufacturing cost assigned.

--Cost of Goods Sold = 1730050 + 1694530 = $ 3,424,580

--Answer = $ 3,424,580